OpenStudy (anonymous):

A flour container is 25cm tall and has a diameter of 20cm. It is filled to 3cm from the top. What volume of flour does it contain? HOw do i solve this please explain!

OpenStudy (mathstudent55):

You need the formula for the volume of a cylinder. V = (pi)(r^2)h, where V = volume of cylinder, pi = famous constant, 3.14159... r = radius of base h = height of cylinder In this case you have the diameter of the base is 20 cm. d = 20 cm r =d/2 = (20 cm)/2 = 10 cm h = 25 cm - 3 cm = 22 cm (you are looking for the volume of the flour, not the total volume of the cylinder. Now insert r = 10 cm and h = 22 cm into V = (pi)(r^2)h, and calculate the volume.

OpenStudy (anonymous):

thank you so much :)

OpenStudy (mathstudent55):

You're welcome.